OpenStudy (anonymous):

In the reaction, CH4 + 2O2 → CO2 + 2H2O, 8.9 × 105 joules of heat are released. Using Einstein’s equation, calculate the mass amount to which this energy equates. [(1 J = 1 ) and the speed of light = 3.00 × 108 ] 9.9 ×1012 kg 9.9 × 10-12 kg 9.9 × 10-14 kg 1.1 × 10-6 kg

OpenStudy (anonymous):

[(1 J = 1 \[\frac{ k*m^2 }{ s^2 }\]) 3.00 × 10^8 \[\frac{ m }{ s }\] ]

OpenStudy (sweetburger):

Wouldn't you just plug in values into E=mc^2?